OpenStudy (inkyvoyd):

e^(ix) = i sin x+cos x. Prove it. Besides using a first order differential equation, and the Maclaurin/Taylor series.

OpenStudy (anonymous):

then i guess it depends on your definiation of these functions. i don't think i have ever seen an explanation without power series, because that is how one usually defines \[e^x\]

OpenStudy (anonymous):

Is this a homework question?

OpenStudy (anonymous):

@inkyvoyd ?

OpenStudy (inkyvoyd):

Nope

OpenStudy (anonymous):

Then follow satellite... probably

OpenStudy (amistre64):

can we use complex numbers? :)

OpenStudy (inkyvoyd):

Sure, just don't go TOO hard.

OpenStudy (inkyvoyd):

"i don't think i have ever seen an explanation without power series, because that is how one usually defines" You can prove that the derivatives of e^(ix) and i sin x+cos x are equivalent, which is one of the 2 ways I know how to prove it.

OpenStudy (amistre64):

i just seen a rather neat limit type proof

OpenStudy (inkyvoyd):

SHOW ME NAO PWEASE

OpenStudy (amistre64):

\[\lim_{n\to inf}(1+(z/n))^n=e^z\] where z = a+bi http://en.wikipedia.org/wiki/Euler's_formula

OpenStudy (inkyvoyd):

HOW DID I MISS THAT

OpenStudy (amistre64):

i just seen it for the first time meself :)

OpenStudy (inkyvoyd):

Bumped from last time; wondering if anyone else has proofs

OpenStudy (amistre64):

i see that :) i just watched a nice youtube presentation on this where he reinvented the series instead of simply applying hte series; but would that be cheating?

OpenStudy (amistre64):

it was something to do with rediscovering the euler formula

OpenStudy (inkyvoyd):

But but but, everyone understands the series!

OpenStudy (amistre64):

the question pretty much reads: prove that 1+1 = 2 without using anything thats been used prior to prove it. :)

OpenStudy (inkyvoyd):

I know, but there are many ways to prove that 1+1 is 2!

OpenStudy (amistre64):

not without reinventing the whell there isnt

OpenStudy (amistre64):

theres another synapse ill never get back lol

OpenStudy (amistre64):

http://www.youtube.com/watch?v=N4ceWhmXxcs i thought this was pretty good, downed a whole bag of microwave popcorn on it last night :)